What is the pKa of the hydrazinium ion, H2N-NH3+? Kb of hydrazine = 8.5*10^-7
Answer #1

Ka = Kw/Kb = (10-14)/(8.5*10-7) = 1.18*10-8

Thus, pKa = -logKa = 7.93